OpenStudy (anonymous):

HELP BEFORE SCHOOL STARTS!! SUMMER WORK! The sum of the angle measures of a pentagon is 540 Degrees. In pentagon ABCDE, angles A, B, and C are congruent, and angle D and E are congruent. The measure of angle D is 30 Degrees greater than the measure of angle A. What is the measure of each angle? Please help i am freaking out, school starts in 20 minutes.

OpenStudy (welshfella):

let measure of angle A be x. Then b = c = x also. Angle d = x + 30 and as e is congruent to d d is also = x + 30.

OpenStudy (welshfella):

so we have the equation 3x + 2(x + 30) = 540 solve this for x

OpenStudy (anonymous):

I think I can do that thank you so much

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

96

OpenStudy (anonymous):

A,B,C=96 D,E=126

OpenStudy (welshfella):

thats correct

OpenStudy (anonymous):

126*2=252 96*3=288 288+252=540

OpenStudy (anonymous):

Answer and proof! Thank you!

OpenStudy (welshfella):

np